OpenStudy (amber15):

Multiply . write a mixed numeral for answer. 70* 7 1/5 = simplify answer my answer is 504 is it right

OpenStudy (fibonaccichick666):

what is \(70 \times 7?\)

OpenStudy (amber15):

490

OpenStudy (fibonaccichick666):

ok now what is \(70 \times \frac{1}{5}\)

OpenStudy (amber15):

14

OpenStudy (amber15):

so the answer is 49.14

OpenStudy (fibonaccichick666):

no you just add them

OpenStudy (fibonaccichick666):

aka yes your answer is correct. A way to think of it is \(70(7+ \frac{1}{5}) \)

OpenStudy (amber15):

oh okay thank you

OpenStudy (fibonaccichick666):

np